Asked the professor and he told me both of #3and #5 are correct. I can understand #3 but may I know why #5is also correct?
Question 4 0 1 point Which of the following conditions must be satisfied in a closed economy? 1) Private saving equals investment. 2) There is no government spending or taxes. 3) Exports equal imports. 4) Government spending equals taxes. 5) Total saving equals investment.
Answer #1

S = I (Total savings=investment)

Here, savings is defined as the funds used for investment.

It can be derived from another identity, which says that in a closed economy with no government, gross domestic product(GDP) equals gross domestic income(GDI):

GDI = C + S

GDP=C + I

Put GDP=GDI

C+S=C+I

Subtract C from both sides to get

S=I

Hence total savings equal investment in a closed economy.